OpenStudy (anonymous):

(sin(theta) + cos(theta))^2 =1 + 2 sin(theta) * cos(theta) prove for the identity

OpenStudy (anonymous):

(sin +cos)^2 = sin^2 +2sin cos + cos^2 = 1+2sin cos (as (sin^2+cos^2)=1)